OpenStudy (hydrogen):

Please help with this simple question! I own a sporting-goods store and must sell ice skates. My budget to buy skates for my store is $4000. I should use half the retail price of the skates as the cost price. The skates I bought cost $149.99 each, and my total came to $3,899.74 (Note: Is this right by the way? Or was I supposed to halve the skates' cost?). They are now for sale in my store. Now, my store has apportioned $2000 as overhead expenses. If my store sold all the skates I purchased, what would be the gross revenue? Gross profit? Net profit? THANK YOU VERY MUCH!!

OpenStudy (mathmale):

The problem states that each pair of skates purchased by the store purchaser cost the store $149.99, and that this cost is HALF the retail price of the pair of skates. How would you go about determining the retail price of each pair? Review the following: Revenue, gross revenue, profit, gross profit, net profit. You'll need a formula for each. Hint: Revenue is the total amount of money taken in by the seller and is equal to the number of pairs of skates sold times the retail selling price. To calculate the net profit, you'd need to calculate the gross revenue (as explained above), and then subtract the seller's cost (of buying these skates wholesale), and, finally, subtract the $2,000 overhead expenses.

OpenStudy (hydrogen):

Thanks for your reply! If the $149.99 cost is half the retail price of the pair of skates, that would mean multiplying x2 and the skates now being sold in the owner's store would equal 299.98 each..? Initially the owner bought 26 pairs of skates for $149.99, totalling to $3,899.74 within the $4k budget If the store sells all the skates that the owner purchased, I came up with a gross revenue of: 3899.74 x 2 = $7,799.48, as the initial skates' cost was half the retail Gross profit: 7799.48 - 3899.74 = $3,899.74 (I feel this is a mistake.. Does that mean the owner did not gain any profit?) Finally, net profit: 7799.48 - 3899.74 - 2000 = $1,899.74 Is this correct at all or have I made some big mistakes..? Thank you very much for your help!!

OpenStudy (mathmale):

"If the $149.99 cost is half the retail price of the pair of skates, that would mean multiplying x2 and the skates now being sold in the owner's store would equal 299.98 each..? " YES, that's correct. We say that the retail (or selling) price of each pair of skates s $299.98. How many pairs of skates can this guy purchase at wholesale, if the max he can spend is $4,000? Divide $4,000 by the wholesale cost of each pair, which is $199.99. Round down your answer to the nearest integer (since you probably weren't planning on selling a fractional pair of skates). Next, figure out the revenue obtained when that many pairs of skates are sold. Mult. $299.98 by the integer you've just calculated. The result is your gross revenue. Subtract $2,000 from that figure, to obtain the net revenue. Doing this is equivalent to finding the net profit: Subtract Costs from Gross Revenue. Any questions about this?

OpenStudy (hydrogen):

4000 ÷ 149.99 = 26.668 -- becoming 26 pairs of skates Gross revenue: 299.98 x 26 = $7,799.48 Net revenue: 7799.48 - 2000 = $5,799.48 Net profit: 5799.48 - 3899.74 = $1,899.74 (Is this correct?) Since by definition gross profit for a retail business is "the amount of revenue left over after the store has paid the costs of the goods that it sold to generate that revenue", does that mean the gross profit is still $3,899.74? (7799.48 - 3899.74?) Thank you very much for all your help and time explaining to me, not sure why I have a lot of trouble with this. Thanks again!!